OpenStudy (anonymous):

Two investments were made totalling $50,000. One investment at 8 1/2% and the other at 6%. If the total earned income was $3500, how much was the investment?

OpenStudy (anonymous):

please answer it, please

OpenStudy (anonymous):

lets put amount invested at 81/2% as x and the rest will therefore be 50,000-x the solve \[.085x=.06(50,000-x)=3500\] to get your answer

OpenStudy (gw2011):

Let x=Investment at 8 1/2% Let y=Investment at 6% 0.085x+0.06y=3000 x+y=3500 Multiply this equation by 0.06 to eliminate y and find x 0.085x+0.06y=3500 0.06x+0.06y=3000 Subtract this equation from the above 0.025x=500 x=$20,000 This is the investment at 8 1/2% y=$30,000 This isthe investment at 6%